# Human Empathy as Encoder: AI-Assisted Depression Assessment in Special Education

### v1

@@ Abstract
Assessing student depression in sensitive environments like special education is challenging. Standardized questionnaires may not fully reflect students’ true situations. Furthermore, automated methods often falter with rich student narratives, lacking the crucial, individualized insights stemming from teachers’ empathetic connections with students. Existing methods often fail to address this ambiguity or effectively integrate educator understanding. To address these limitations by fostering a synergistic human-AI collaboration, this paper introduces Human Empathy as Encoder (HEAE), a novel, human-centered AI framework for transparent and socially responsible depression severity assessment. Our approach uniquely integrates student narrative text with a teacher-derived, 9-dimensional "Empathy Vector" (EV), its dimensions guided by the PHQ-9 framework,to explicitly translate tacit empathetic insight into a structured AI input enhancing rather than replacing human judgment. Rigorous experiments optimized the multimodal fusion, text representation, and classification architecture, achieving 82.74% accuracy for 7-level severity classification. This work demonstrates a path toward more responsible and ethical affective computing by structurally embedding human empathy.
@@
Human-AI Collaboration, Depression Assessment ,Empathy Encoding, AI in Special Education
@@ I Introduction
The motivation for this research stems from firsthand observations in special education settings, where supporting student mental well-being, particularly regarding depression, is paramount. These environments, with their typically small student groups, foster deep teacher familiarity with individual needs and contexts. However, depression assessment here presents unique challenges: standardized self-report questionnaires like PHQ-9 [1] often fail to capture students’ authentic emotional states, while their natural conversations and writings can reveal richer insights. This creates a critical tension—educators possess invaluable empathy and contextual understanding from sustained interactions, yet without extensive psychological training, they struggle to systematically interpret vague narrative texts for signs of depression.
The challenge of interpreting student narratives is further compounded when considering purely automated approaches. Existing affective computing models often struggle with the nuances of expressed mental states like depression, potentially oversimplifying the assessment [2]. Moreover, the growing adoption of large language models, especially cloud-based services, raises significant privacy concerns regarding sensitive student data [3]. Thus, the limitations of standalone human interpretation (highlighted previously) and the shortcomings of current AI tools collectively demand a new generation of assessment solutions grounded in social responsibility and synergistic human-AI collaboration [4]
To navigate these complex challenges, this paper introduces the "Human Empathy as Encoder (HEAE)" framework, a novel human-centered AI paradigm. HEAE fundamentally reorients depression assessment by structurally integrating teachers’ tacit empathetic insights and experiential understanding directly into the AI analysis process. This approach complements and amplifies human judgment rather than attempting to replace it, fosters greater transparency compared to opaque automated systems and captures individualized nuances often missed by standardized questionnaires. Our work addresses implementation challenges through an innovative LLM-assisted annotation pipeline leveraging "golden seeds"—expert-annotated examples, to generate high-quality, nuanced training data for the HEAE model. Architecturally, HEAE embodies the principle of "structured simplicity + targeted complexity" - a design philosophy that strategically allocates model sophistication where it most benefits human-AI collaboration. Beyond achieving 82.74% accuracy in 7-level depression severity classification, this work provides both a privacy-preserving assessment tool for educational settings and a broader template for developing responsible, human-centered affective computing systems.
@@ III-B1 Golden Seed Development
To establish a high-quality foundation and encapsulate our target annotation logic, we created 200 initial examples termed "golden seeds." Raw narratives were sourced from Reddit forums, with all user-identifying information removed.
These golden seeds underwent manual annotation by experts with substantial experience in special education settings. This crucial step defined two key components. First, we established a machine-interpretable mapping between narrative content and scores across nine EV dimensions using the previously described 0-5 severity scale. Second, we applied 7-level classification criteria (outlined in the Appendix) to convert these dimensional scores into the final severity classification. Each golden seed included the narrative text, the 9-dimensional EV, final severity label, and detailed explanations justifying the scores, thus encapsulating the target reasoning process.
@@ III-B2 Novel Automated Annotation Pipeline
To efficiently scale our dataset while maintaining consistent interpretation patterns, we developed an LLM-assisted annotation pipeline (Figure 1). This approach leverages 20 carefully selected golden seeds as few-shot demonstrations for the LLM, To further guide the LLM and ensure annotation consistency, the prompt was supplemented with explicit details regarding the Empathy Vector (EV) structure and the 7-level classification criteria (as outlined in the Appendix), alongside instructions for the LLM to adopt a persona reflecting expertise in empathetic mental well-being assessment in educational contexts.
@@ III-B4 Model Inputs, Target
The model utilizes two primary inputs: (i) Student Narrative Text, and (ii) EV (9-dim, 0-5 scale from teacher interpretation). The classification target is the 7-level depression severity assigned during data generation.
@@ III-C Final Model Architecture
Figure 2: The final architecture of the Human Empathy as Encoder model, detailing text processing, EV processing, asymmetric cross-modal enhancement, and the deep classifier head"
@@ III-D Experimental Setup and Training
Our curated dataset (detailed in Section III-B3) was partitioned into 90% for training and 10% for validation. We employed the AdamW optimizer with a learning rate of 1e-5 and a batch size of 8. Model performance was evaluated using accuracy and macro F1-score on the validation set.
@@ IV-B Optimizing Multimodal Fusion
IV-B1 Evaluating Fusion Strategies: From Standard Attention to Novel Enhancements
To effectively integrate the text and EV modalities, we first investigated the performance of standard multi-head cross-attention mechanisms, a common approach for multimodal fusion. As shown in Figure 3, even the most effective configuration of these attention models offered only a slight improvement over the simple concatenation baseline. This modest outcome motivated our exploration of more tailored fusion strategies. Consequently, we developed and evaluated our advanced fusion mechanisms: Symmetric Cross-Modal Enhancement (SCME) and Asymmetric Cross-Modal Enhancement (ACME) (with ACME’s design detailed in Section III-C4). In stark contrast, while SCME demonstrated further improvement, our ACME approach achieved significantly superior performance, reaching 77.12% accuracy. This result clearly outperformed both the baseline and all tested attention-based models (see Figure 3), thus validating this new direction.
Figure 3: Comparison of different multimodal fusion methods. Our optimized Asymmetric Cross-Modal Enhancement (ACME) significantly outperforms Attention-Based methods and Symmetric Cross-Modal Enhancement (SCME).
@@ IV-B2 Optimizing Asymmetric Cross-Modal Enhancement
Based on this observation, we tested fixed asymmetric factors within the ACME configuration. Prioritizing the EV’s influence on text (setting $\alpha_{e\rightarrow t}=0.85$) while reducing the reverse influence (setting $\alpha_{t\rightarrow e}=0.30$) yielded the best performance in this series, reaching 77.12% accuracy (Figure 4).
Figure 4: Optimization of asymmetric enhancement factors for the ACME mechanism
@@ IV-C Text Representation Optimization
Evaluating standard document pooling strategies (Figure 5), Max Pooling demonstrated clearly superior performance over alternatives like Mean Pooling, Attention Pooling, and hybrid approaches. This highlighted the importance of capturing peak emotional signals from the narrative texts for assessing depression severity. Further significant improvement was achieved by incorporating Layer Normalization after Max Pooling, yielding an accuracy of 79.08% and establishing ’Max Pooling + LayerNorm’ as our optimal text representation strategy.
Importantly, while Max Pooling identifies these "peak signals," these are not interpreted naively (e.g., based solely on keywords such as "I want to die") within the HEAE framework. Instead, the asymmetric fusion mechanism (Sec IV-B2) leverages the teacher-derived EV to provide essential context. This effectively grounds and validates text-based signals against the teacher’s holistic, empathetic assessment, underscoring the indispensable role of structured human empathy in achieving nuanced and reliable judgment from student narratives.
@@ IV-D Classifier Head Optimization
With optimized fusion (Sec IV-B2) and text representation (Sec IV-C), we refined the classifier head to better handle the fine-grained 7-level severity classification. We found that increasing the classifier depth by implementing a multi-layer architecture ((896D → 448D → 224D → 7 classes) significantly improved performance. Within this deeper structure, GELU activation yielded the best performance, outperforming ReLU and SiLU. Further enhancement was achieved by incorporating GLUs into this GELU-activated structure, bringing the final accuracy to 82.74% (Figure 6).
This classifier’s improved performance highlights a key architectural principle: unlike earlier model stages (e.g., fusion, text pooling) where simpler mechanisms were optimal for preserving key signals, the final classification of rich, aggregated multimodal features demands significantly increased representational capacity. The deeper, gated architecture, incorporating GELU activation and GLUs, provided this necessary capacity, enabling effective learning of complex non-linearities and adaptive features to discern subtle differences among the 7 severity levels from the consolidated inputs.
Figure 6: Comparison of different classifier architectures, showing that the combination of GELU activation with GlUs achieves the best performance.
@@ IV-E Final Model Performance
The final optimized HEAE model, incorporating all optimal components (asymmetric fusion, max pooling, and deep GELU+GLUs classifier head), achieved 82.74% accuracy and 82.77% macro F1-score on the validation set for our 7-level depression severity classification task. This represents a significant improvement of approximately 10 percentage points over the baseline model (73% accuracy), demonstrating the effectiveness of our human-empathy guided approach.
@@ V-A Responsible Design Principles for Human-AI Collaboration
Our experiments reveal a key design principle for human-AI collaboration: complexity should be strategically allocated, not uniformly applied. The relative simplicity in modal fusion and representation extraction contrasts sharply with the targeted depth in classification. This reflects more than just technical choices; it embodies a philosophy regarding the transparency of the human-AI interaction. This "structured simplicity + targeted complexity" approach challenges the "more complex is better" tendency often seen in the pursuit of model performance. It suggests that truly effective human-AI systems should respect and leverage the respective strengths of both humans and AI. Therefore, we advocate prioritizing understandability and signal fidelity during early-stage information integration. At the same time, necessary computational sophistication should be permitted for the final, fine-grained decision-making stage. This balancing principle extends beyond specific algorithm choices and aligns with current ethical demands for AI transparency and explainability. It charts a new course for designing future responsible affective computing systems—ones that are both effective and ethically considerate, especially in highly sensitive areas like special education
@@ Acknowledgments
This work is dedicated to the one who taught me the meaning of pure love and belief — a soul whose quiet presence still guides this research.

### v2

@@ Abstract
Assessing student depression in sensitive environments like special education is challenging. Standardized questionnaires may not fully reflect students’ true situations. Furthermore, automated methods often falter with rich student narratives, lacking the crucial, individualized insights stemming from teachers’ empathetic connections with students. Existing methods often fail to address this ambiguity or effectively integrate educator understanding. To address these limitations by fostering a synergistic human-AI collaboration, this paper introduces Human Empathy as Encoder (HEAE), a novel, human-centered AI framework for transparent and socially responsible depression severity assessment. Our approach uniquely integrates student narrative text with a teacher-derived, 9-dimensional "Empathy Vector" (EV), its dimensions guided by the PHQ-9 framework,to explicitly translate tacit empathetic insight into a structured AI input enhancing rather than replacing human judgment. Rigorous experiments optimized the multimodal fusion, text representation, and classification architecture, achieving 82.05% ± 0.58% accuracy via 5-fold cross-validation for 7-level severity classification. This work demonstrates a path toward more responsible and ethical affective computing by structurally embedding human empathy.
@@
Human-AI Collaboration, Depression Assessment, Empathy Encoding, Responsible AI
@@ I Introduction
The motivation for this research stems from firsthand observations in special education settings, where supporting student mental well-being, particularly regarding depression, is paramount. These environments, with their typically low student-to-teacher ratios, foster deep teacher familiarity with individual needs and contexts. However, depression assessment here presents unique challenges: standardized self-report questionnaires like PHQ-9 [1] often fail to capture students’ authentic emotional states, while their natural conversations and writings can reveal richer insights. This creates a critical tension—educators possess invaluable empathy and contextual understanding from sustained interactions, yet without extensive psychological training, they struggle to systematically interpret vague narrative texts for signs of depression.
The challenge of interpreting student narratives is further compounded when considering purely automated approaches. Existing affective computing models often struggle with the nuances of expressed mental states like depression, potentially oversimplifying the assessment [2]. Moreover, the growing adoption of large language models, especially cloud-based services, raises significant privacy concerns regarding sensitive student data [3]. Thus, the limitations of standalone human interpretation and the shortcomings of current AI tools collectively demand a new generation of assessment solutions grounded in social responsibility and synergistic human-AI collaboration [4]
To navigate these complex challenges, this paper introduces the "Human Empathy as Encoder (HEAE)" framework, a novel human-centered AI paradigm. HEAE fundamentally reorients depression assessment by structurally integrating teachers’ tacit empathetic insights and experiential understanding directly into the AI analysis process. This approach complements and amplifies human judgment rather than attempting to replace it, fosters greater transparency compared to opaque automated systems and captures individualized nuances often missed by standardized questionnaires. Our work addresses implementation challenges through an innovative LLM-assisted annotation pipeline leveraging "golden seeds"—expert-annotated examples, to produce high-quality, nuanced training data for the HEAE model. Architecturally, HEAE embodies the design principle of "structured simplicity + targeted complexity," which strategically allocates model sophistication where it most benefits human-AI collaboration. Beyond achieving 82.05% ± 0.58% accuracy via 5-fold cross-validation accuracy in 7-level depression severity classification, this work provides both a privacy-preserving assessment tool for educational settings and a broader template for developing responsible, human-centered affective computing systems.
@@ III-B1 Golden Seed Development
To establish a high-quality foundation and encapsulate our target annotation logic, we created a set of initial examples termed "golden seeds." Raw narratives were sourced from Reddit forums, with all user-identifying information removed.
These golden seeds underwent manual annotation by experts with substantial experience in special education settings. This crucial step defined two key components.First, we established a machine-interpretable mapping between narrative content and scores across nine EV dimensions using the previously described 0-5 severity scale for each dimension, without calculating a total score. Second, these expert-generated EV scores were converted into a final 7-level depression severity classification. To ensure this EV-to-severity conversion method aligns with established understanding, its consistency with the PHQ-9’s approach to categorizing severity was validated on all 200 golden seeds, yielding a Cohen’s Kappa of 0.705, indicating substantial inter-rater agreement, and an overall accuracy of 75.5%. Each golden seed subsequently included the narrative text, its expert-generated 9-dimensional EV, the resulting final 7-level severity label, and detailed explanations justifying the EV scores, thus encapsulating the target reasoning process.
@@ III-B2 Novel Automated Annotation Pipeline
To efficiently scale our dataset while maintaining consistent interpretation patterns, we developed an LLM-assisted annotation pipeline (Figure 1). From the full set of 200 validated golden seeds, we carefully selected 20 diverse examples to serve as few-shot demonstrations for the LLM. This number was chosen to provide a rich variety of narrative styles and severity levels, while respecting the input token limitations of the language model. To further guide the LLM and ensure annotation consistency, the prompt was supplemented with explicit details regarding the Empathy Vector (EV) structure and our validated EV-to-severity conversion method, alongside instructions for the LLM to adopt a persona reflecting expertise in empathetic mental well-being assessment in educational contexts.
@@ III-B4 Model Inputs, Target
The model utilizes two primary inputs: (i) Narrative Text, and (ii) EV (9-dim, 0-5 scale from teacher interpretation). The classification target is the 7-level depression severity assigned during data generation.
@@ III-C Final Model Architecture
Figure 2: The architecture of the proposed Human Empathy as Encoder (HEAE) model.
@@ III-D Experimental Setup and Training
Our curated dataset (detailed in Section III-B3) was evaluated using 5-fold cross-validation to ensure robust performance assessment. We employed the AdamW optimizer with a learning rate of 1e-5 and a batch size of 8. Model performance was evaluated using accuracy and macro F1-score across all folds.
@@ IV-B Optimizing Multimodal Fusion
IV-B1 Evaluating Fusion Strategies
To effectively integrate the text and EV modalities, we first investigated the performance of standard multi-head cross-attention mechanisms, a common approach for multimodal fusion. As shown in Figure 3, even the most effective configuration of these attention models offered only a slight improvement over the simple concatenation baseline. This modest outcome motivated our exploration of more tailored fusion strategies. Consequently, we developed and evaluated our advanced fusion mechanisms: Symmetric Cross-Modal Enhancement (SCME) and Asymmetric Cross-Modal Enhancement (ACME). The design of the ACME mechanism is detailed in Section III-C4. In stark contrast, while SCME demonstrated further improvement, our ACME approach achieved significantly superior performance, reaching 77.12% accuracy. This result clearly outperformed both the baseline and all tested attention-based models, thus validating this new direction.
Figure 3: Performance comparison of different multimodal fusion strategies.
@@ IV-B2 Optimizing Asymmetric Cross-Modal Enhancement
Based on this observation, we tested several fixed asymmetric factor configurations to identify the optimal setup. As shown in Table I, the configuration with $\alpha_{e\rightarrow t}=0.85$ and $\alpha_{t\rightarrow e}=0.30$ achieved the best performance at 77.12% accuracy, outperforming both other asymmetric variants and dynamic factor learning.
TABLE I: Performance of ACME with Different Asymmetric Factor Configurations.
Configuration Factors ($\alpha_{e\rightarrow t}$ / $\alpha_{t\rightarrow e}$) Accuracy (%)
Asymmetric Variant 1 0.90 / 0.25 76.54
ACME (Optimal) 0.85 / 0.30 77.12
Asymmetric Variant 2 0.80 / 0.35 76.16
Dynamic Factor Learned 76.79
@@ IV-C Text Representation Optimization
Evaluating standard document pooling strategies (Figure 4), Max Pooling demonstrated clearly superior performance over alternatives like Mean Pooling, Attention Pooling, and Adaptive Hierarchical pooling. This highlighted the importance of capturing peak emotional signals from the narrative texts for assessing depression severity. Further significant improvement was achieved by incorporating Layer Normalization after Max Pooling, yielding an accuracy of 79.08% and establishing ’Max Pooling + LayerNorm’ as our optimal text representation strategy.
Importantly, while Max Pooling identifies these "peak signals," these are not interpreted naively (e.g., based solely on keywords such as "I want to die") within the HEAE framework. Instead, the ACME mechanism (Sec IV-B2) leverages the teacher-derived EV to provide essential context. This effectively grounds and validates text-based signals against the teacher’s holistic, empathetic assessment, underscoring the indispensable role of structured human empathy in achieving nuanced and reliable judgment from student narratives.
@@ IV-D Classifier Head Optimization
With optimized fusion (Sec IV-B2) and text representation (Sec IV-C), we refined the classifier head to better handle the fine-grained 7-level severity classification. We found that increasing the classifier depth by implementing a multi-layer architecture ((896D → 448D → 224D → 7 classes) significantly improved performance. Within this deeper structure, GELU activation yielded the best performance, outperforming ReLU and SiLU. Further enhancement was achieved by incorporating GLUs into this GELU-activated structure, bringing the final accuracy to 82.05% (Figure 5).
This classifier’s improved performance highlights a key architectural principle: unlike earlier model stages (e.g., fusion, document pooling) where tailored simpler mechanisms were optimal for preserving key signals, the final classification of rich, aggregated multimodal features demands significantly increased representational capacity. The deeper, gated architecture, incorporating GELU activation and GLUs, provided this necessary capacity, enabling effective learning of complex non-linearities and adaptive features to discern subtle differences among the 7 severity levels from the consolidated inputs.
Figure 5: Performance comparison of different classifier head architectures.
@@ IV-E Final Model Performance
The final optimized HEAE model, incorporating all optimal components (asymmetric fusion, max pooling, and deep GELU+GLUs classifier head), achieved 82.05% ± 0.58% accuracy and 82.08% macro F1-score for our 7-level depression severity classification task. This represents a significant improvement of approximately 9% points over the baseline model (73% accuracy), demonstrating the effectiveness of our human-empathy guided approach.
@@ V-A Responsible Design Principles for Human-AI Collaboration
Our experiments reveal a key design principle for human-AI collaboration: complexity should be strategically allocated, not uniformly applied. The tailored simplicity in modal fusion and representation extraction contrasts sharply with the targeted depth in classification. This reflects more than just technical choices; it embodies a philosophy regarding the transparency of the human-AI interaction. This "structured simplicity + targeted complexity" approach challenges the "more complex is better" tendency often seen in the pursuit of model performance. It suggests that truly effective human-AI systems should respect and leverage the respective strengths of both humans and AI. Therefore, we advocate prioritizing understandability and signal fidelity during early-stage information integration. At the same time, necessary computational sophistication should be permitted for the final, fine-grained decision-making stage. This balancing principle extends beyond specific algorithm choices and aligns with current ethical demands for AI transparency and explainability. It charts a new course for designing future responsible affective computing systems—ones that are both effective and ethically considerate, especially in highly sensitive areas like special education
@@ VII Conclusion
Data Availability Statement
The "Golden Seeds" dataset, validation scripts, and the source code for our HEAE model implementation will be made publicly available in an online repository upon publication of this article. The link will be provided in the final camera-ready version.
@@ Acknowledgments
This work is dedicated to the one whose profound inspiration and unwavering belief have reshaped my life -a soul whose quiet presence still guides this research.